OpenStudy (anonymous):

A painter used 5/8 gallon of paint for one job and 1/4 gallon for another job. How many gallons of paint did the painter use for both jobs?

OpenStudy (anonymous):

1st job= 5/8 2nd job= 1/4=2/8 Add. 5/8 + 2/8 = 7/8 Gallons used for both jobs.